# Disentangling trust from cooperation: Trust as reduced monitoring across social dilemmas

### v3

@@ Abstract
Keywords: Trust, Costly monitoring, Decision heuristic, Evolutionary game theory, Social dilemmas.
@@ 1 Introduction
Trust underlies cooperation in every human society [1, 2], and hence has been called “the glue of healthy societies and the grease of economic productivity” [3]. This implies that the Artificial Intelligence (AI) agents of the future will also need the capability to build trust, in order to cooperate effectively both with each other and with people [4, 5]. However, this raises a fundamental question: what is the individual adaptive benefit of trusting another agent? By trusting another, an agent exposes itself to risk [6] – the risk that the other agent might not act in an honest and reliable way.
This question has been difficult to answer because there are many different verbal definitions of trust in both social science and artificial intelligence. These definitions can be categorised into: 1. behavioural definitions, which equate trust with more cooperative choices in a game, and 2. psychological definitions, which equate trust with psychological states such as viewing another individual as benevolent and acting with integrity [7]. Prior game-theoretic work on trust in economics [8] and multi-agent systems [9] has typically followed 1., and effectively equated trust with high levels of cooperation. The risky investment decision modelled by the “trust game” [10, 11, 12], which is effectively a one-sided Prisoner’s Dilemma [13], provides a paradigm example. The fundamental problem with this is that it conflates an outcome, cooperation, with one of its causes, trust [14, 15, 16]. As such, these definitions of trust do not allow us to delineate the role of trust in promoting cooperation. An alternative tradition in multi-agent systems has followed 2., and equipped agents with cognitive mechanisms to make trust decisions [17]. However, this has still not resulted in a general and mechanistic way to measure trust that applies in a wide range of settings. This is because it defines trust relative to an agent’s internal state, which may not be observable in an interpretable way, e.g. as is the case for current Large Language Model (LLM) agents, as well as for human agents.
@@
Trust as reduced monitoring fits the intuitive folk psychology rule of “If I trust you then I don’t have to keep checking what you’re doing”. This is in accordance with theories from social science; in particular, Luhmann argues that human agents use a trust heuristic to avoid having to calculate the likelihood of every possible outcome of an interaction [6, 7]. When trust is established, the adaptive benefit is that people save the time and effort of trying to exhaustively calculate whether or not they should cooperate and take the risk of being exploited. This has often been hinted at in the definitions of trust used in artificial intelligence [18]. For example, in Taddeo’s influential definition of e-trust, greater levels of trust entail less effort spent supervising the trusted agent [19]. Similarly, in Castelfranchi and Falcone’s notion of strict trust, trust is antagonistic to monitoring, where monitoring is defined as an action “aimed at ascertaining whether another action has been successfully executed or if a given state of the world has been realized or maintained…” [20, pg. 193]. Trust as a decision to reduce resources (time and effort) invested into monitoring has also been proposed as a pragmatic definition of trust between people and artificial intelligence [21, 22, 23, 24, 25].
@@
In repeated games, agents interact with each other several times (rounds). This gives agents the possibility to condition their behaviour on how the other agent has acted previously, for example, by using reciprocal strategies such as “cooperate this time if the other agent cooperated on the previous round”. These reciprocal strategies – where agents can base their knowledge of how the other agent has acted previously on either their own first-hand knowledge (direct reciprocity) or information from third parties (indirect reciprocity) – underpin cooperation in human societies [26]. But crucially, to be able to use reciprocal strategies, an agent needs to know whether the observed actions of another agent are actually cooperative or not. It is traditionally assumed that once an agent observes an action, it can instantly and with zero cost determine whether that observed action was cooperative. In other words, monitoring the actions of a co-player is costless. However, in reality, monitoring whether an observed action was actually cooperative carries an opportunity cost, representing the time, effort and resources spent. For example, in e-commerce it may not be easy for a buyer to tell whether a seller sent them goods of the advertised quality, rather than say a fake, and so choosing to monitor this may take a substantial amount of their time. Similarly, monitoring the actions of AI agents may involve simulating execution of their source code [27], the cost of which would depend on the power of the simulating agent and the auditing frameworks in place. Thus, in most real-world settings, conditional strategies such as Tit-for-Tat would need to pay a monitoring cost in order to know if their co-player cooperated on the previous round or not.
@@ 2.2.1 Trust-based strategies
These can be understood as follows. In the first $\theta$ rounds, before the trust threshold is reached, both TUC and TUD play C and keep monitoring their partners’ actions. In each of these rounds they therefore both obtain a payoff $R-\varepsilon$. Once the trust threshold $\theta$ is reached, TUC will monitor the action of its partner only occasionally with probability $p$, while TUD will switch to defecting. Suppose that in the next round TUC verifies its partner’s action and so finds them to be defecting. In this case, it obtains $S$ in the current round and $P-\varepsilon$ in the remaining rounds, since it reverts to TFT. Otherwise, i.e. if TUC does not monitor in that round (with probability 1p1-p), the process above is iterated for the payoffs calculation.
@@ 3.3 Trust-as-reduced-monitoring when individuals make errors
4 Discussion
Our results can be summarised in three points. First, we have shown that trust, as a heuristic to avoid the cost of always monitoring a partner’s actions, is a successful strategy that can evolve in a wide range of interactions. Importantly, this happens even in the presence of exploitative strategies (TUD) tailored to take advantage of this trust.
@@ 4 Discussion
Our operationalisation of trust – as choosing not to monitor another agent’s actions – provides an objective behavioural measure that applies across a broad spectrum of social interactions. Using this, trust can be measured in experimental settings by how often someone chooses (or not) to pay a cost to monitor another agent’s actions. This could be carried out in standard behavioural economics settings [48], where a participant is given the option to pay to find out the action of its partner in each round of a repeated social dilemma game. Some experimental studies have already considered costly monitoring for the particular case of the “trust game” [49, 50]. In this game, one player is the investor and the other player is the trustee. The investor begins with an endowment amount of money. They then choose a proportion (between 0 and 1 inclusive) of this to send to the trustee. The amount sent is multiplied by a factor >1>1. The trustee then chooses the proportion of this to return to the investor. When the game is repeated, most experiments assume that the investor is given complete information about the previous actions of the trustee at no cost, i.e. that monitoring is costless. However, a study has shown that when a cost is introduced to obtain this information, people are sensitive to this cost and monitor less [49]. Moreover, the frequency of monitoring was lower in later rounds, in a pattern akin to the TUC strategy. Future work should extend these experiments beyond the trust game and into the larger space of social dilemma games that we have modelled.
Crucially, this trust measure can also be applied in more realistic experimental settings. One important application of this is experiments to measure the factors affecting people’s trust in AI. Although there have been many empirical studies in computer science that have aimed to measure factors affecting trust in various applications of AI, these studies have lacked a unified theoretical framework [51, 52]. This has led to conflicting results and a lack of predictive models, with “Trustworthiness mechanisms and measures […] being advanced in AI regulations and standards that may not actually increase trust” [53]. The trust measure we have introduced here can help to resolve this impasse, by providing an objective measure of trust that is grounded in (evolutionary) game theory. For example, incentivised experiments could be designed where a user can take advice from an AI agent in a realistic setting, such as an e-commerce chatbot recommending products or an LLM answering questions about a text, and can pay a cost to verify (monitor) whether that advice is correct. This can then be connected with EGT models by varying factors such as how frequently the AI gives correct information (i.e. its strategy to cooperate or defect), or whether the AI has a high reputation from other users.
@@
In conclusion, we have developed a behavioural measure of trust that applies across the spectrum of symmetric two-player social dilemma games, and that distinguishes observed trust from observed cooperation. Because this behavioural measure does not depend on access to an agent’s internal state, it can apply to both human and AI agents. Using the methods of evolutionary game theory, we have shown that Trust-based Cooperation, which verifies its partners’ actions only occasionally after a trust-threshold is reached, is a successful strategy in a wide range of social interactions. We have also shown that the presence of trust increases the overall level of cooperation in a population, especially when agents can make unintentional errors in their actions. This adds formal support for the idea that high-trust societies lead to more cooperative and friction-free economies [1].
@@
The authors used AI-based tools (ChatGPT 5.1 and Elicit) solely to assist with the identification of potentially relevant literature. All cited sources were independently reviewed and evaluated by the authors. AI tools were not used for data analysis, interpretation, or generation of manuscript text.

### v4

# Disentangling trust from cooperation: Evolution of trust as reduced monitoring in social dilemmas

@@ Abstract
Keywords: Trust, Cooperation, Costly monitoring, Social dilemmas, Evolutionary game theory.
@@ 1 Introduction
Trust underlies cooperation in every human society [1, 2], and hence has been called “the glue of healthy societies and the grease of economic productivity” [3]. This implies that the Artificial Intelligence (AI) agents of the future will also need the capability to build trust, to effectively cooperate [4, 5, 6] with each other and with humans [7, 8]. However, this raises a fundamental question: what is the individual adaptive benefit of trusting another agent? By trusting another, an agent exposes itself to risk [9] – the risk that the other agent might not act in an honest and reliable way.
This question has been difficult to answer because there are many different verbal definitions of trust in both social science and artificial intelligence. These definitions can be categorised into: 1. behavioural definitions, which equate trust with more cooperative choices in a game, and 2. psychological definitions, which equate trust with psychological states such as viewing another individual as benevolent and acting with integrity [10]. Prior game-theoretic work on trust in economics [11] and multi-agent systems [12] has typically followed 1., and effectively equated trust with high levels of cooperation. The risky investment decision modelled by the Trust Game [13, 14, 15], which is effectively a one-sided Prisoner’s Dilemma [16], provides a paradigm example. The fundamental problem with this is that it conflates an outcome, cooperation, with one of its causes, trust [17, 18, 19]. As such, these definitions of trust do not allow us to delineate the role of trust in promoting cooperation. An alternative tradition in multi-agent systems has followed 2., and equipped agents with cognitive mechanisms to make trust decisions [20]. However, this has still not resulted in a general and mechanistic way to measure trust that applies in a wide range of settings. This is because it defines trust relative to an agent’s internal state, which may not be observable in an interpretable way, e.g. as is the case for current Large Language Model (LLM) agents, as well as for human agents.
@@
Trust as reduced monitoring fits the intuitive folk psychology rule of “If I trust you then I don’t have to keep checking what you’re doing”. This is in accordance with theories from social science; in particular, Luhmann argues that human agents use a trust heuristic to avoid having to calculate the likelihood of every possible outcome of an interaction [9, 10]. When trust is established, the adaptive benefit is that people save the time and effort of trying to exhaustively calculate whether or not they should cooperate and take the risk of being exploited. This has often been hinted at in the definitions of trust used in artificial intelligence [21]. For example, in Taddeo’s influential definition of e-trust, greater levels of trust entail less effort spent supervising the trusted agent [22]. Similarly, in Castelfranchi and Falcone’s notion of strict trust, trust is antagonistic to monitoring, where monitoring is defined as an action “aimed at ascertaining whether another action has been successfully executed or if a given state of the world has been realized or maintained…” [23, pg. 193]. Trust as a decision to reduce resources (time and effort) invested into monitoring has also been proposed as a pragmatic definition of trust between people and artificial intelligence [24, 25, 26, 27, 28].
@@
In repeated games, agents interact with each other several times (rounds). This gives agents the possibility to condition their behaviour on how the other agent has acted previously, for example, by using reciprocal strategies such as “cooperate this time if the other agent cooperated on the previous round”. These reciprocal strategies – where agents can base their knowledge of how the other agent has acted previously on either their own first-hand knowledge (direct reciprocity) or information from third parties (indirect reciprocity) [29] – underpin cooperation in human societies [30]. But crucially, to be able to use reciprocal strategies, an agent needs to know whether the observed actions of another agent are actually cooperative or not. It is traditionally assumed that once an agent observes an action, it can instantly and with zero cost determine whether that observed action was cooperative. In other words, monitoring the actions of a co-player is costless.
However, in reality, monitoring whether an observed action was actually cooperative carries an opportunity cost, representing the time, effort and resources spent that could be allocated to other productive activities. For example, in e-commerce it may not be easy for a buyer to tell whether a seller sent them goods of the advertised quality, rather than say a fake, and so choosing to monitor this may take a substantial amount of their time. Although reputational information can sometimes reduce the need for this kind of direct monitoring (see, e.g., reputation mechanisms in Trust Games [31, 32]), reputational information itself can need verification for accuracy. For instance, determining whether reviews about a seller or service provider are genuine or fake can be difficult and time-consuming, and these costs may increase in the future given the ability of LLMs to automate the generation of plausible-sounding review text [33, 34]. Finally, monitoring the actions of AI agents can also be very costly, as this may involve simulating execution of their source code [35], the cost of which would depend on the power of the simulating agent and the auditing frameworks in place. Thus, in most real-world settings, conditional strategies such as Tit-for-Tat would need to pay a monitoring cost in order to know if their co-player cooperated on the previous round or not.
@@ 2.2.1 Trust-based strategies
These can be understood as follows. In the first $\theta$ rounds, before the trust threshold is reached, both TUC and TUD play C and keep monitoring their partners’ actions. In each of these rounds they therefore both obtain a payoff $R-\varepsilon$. Once the trust threshold $\theta$ is reached, TUC will monitor the action of its partner only occasionally with probability $p$, while TUD will switch to defecting. Suppose that in the next round TUC monitors its partner’s action and so finds them to be defecting. In this case, it obtains $S$ in the current round and $P-\varepsilon$ in the remaining rounds, since it reverts to TFT. Otherwise, i.e. if TUC does not monitor in that round (with probability 1p1-p), the process above is iterated for the payoffs calculation.
@@ 3.3 Trust-as-reduced-monitoring when individuals make errors
4 Conclusion and implications
Our results can be summarised by three points. First, we have shown that trust, as a heuristic to avoid the cost of always monitoring a partner’s actions, is a successful strategy that can evolve in a wide range of interactions. Importantly, this happens even in the presence of exploitative strategies (TUD) tailored to take advantage of this trust.
@@ 4 Conclusion and implications
Our operationalisation of trust – as choosing not to monitor another agent’s actions – provides an objective behavioural measure that applies across a broad spectrum of social interactions. Using this, trust can be measured in experimental settings by how often someone chooses (or not) to pay a cost to monitor another agent’s actions. This could be carried out in standard behavioural economics settings [56], where a participant is given the option to pay to find out the action of its partner in each round of a repeated social dilemma game. Some experimental studies have already considered costly monitoring for the particular case of the Trust Game [57, 58]. In this game, one player is the investor and the other player is the trustee. The investor begins with an endowment amount of money. They then choose a proportion (between 0 and 1 inclusive) of this to send to the trustee. The amount sent is multiplied by a factor >1>1. The trustee then chooses the proportion of this to return to the investor. When the game is repeated, most experiments assume that the investor is given complete information about the previous actions of the trustee at no cost, i.e. that monitoring is costless. However, a study has shown that when a cost is introduced to obtain this information, people are sensitive to this cost and monitor less [57]. Moreover, the frequency of monitoring was lower in later rounds, in a pattern akin to the TUC strategy. Future work should extend these experiments beyond the trust game and into the larger space of social dilemma games that we have modelled.
Crucially, this trust measure can also be applied in more realistic experimental settings. One important application of this is experiments to measure the factors that affect people’s trust in AI. Although there have been many empirical studies in computer science that have aimed to measure factors affecting trust in various applications of AI, these studies have lacked a unified theoretical framework [59, 60]. This has led to conflicting results and a lack of predictive models, with “Trustworthiness mechanisms and measures […] being advanced in AI regulations and standards that may not actually increase trust” [61]. The trust measure we have introduced here can help to resolve this impasse, by providing an objective measure of trust that is grounded in (evolutionary) game theory. For example, incentivised experiments could be designed where a user can take advice from an AI agent in a realistic setting, such as an e-commerce chatbot recommending products or an LLM answering questions about a text, and can pay a cost to verify (monitor) whether that advice is correct. This can then be connected with EGT models by varying factors such as how frequently the AI gives correct information (i.e. its strategy to cooperate or defect), or whether the AI has a high reputation from other users.
@@
In conclusion, we have developed a behavioural measure of trust that applies across the spectrum of symmetric two-player social dilemma games, and that distinguishes observed trust from observed cooperation. Because this behavioural measure does not depend on access to an agent’s internal state, it can apply to both human and AI agents. Using the methods of evolutionary game theory, we have shown that Trust-based Cooperation, which monitors its partners’ actions only occasionally after a trust-threshold is reached, is a successful strategy in a wide range of social interactions. We have also shown that the presence of trust increases the overall level of cooperation in a population, especially when agents can make unintentional errors in their actions. This adds formal support for the idea that high-trust societies lead to more cooperative and friction-free economies [1].
@@
During the preparation of this work the authors used AI-based tools (ChatGPT with Deep Research and Elicit) to assist with the identification of potentially relevant literature. All cited sources were independently read, evaluated and integrated into the manuscript by the authors. The authors take full responsibility for the content of the publication.